# Large-Scale Continual Scheduling and Execution for Dynamic Distributed Satellite Constellation Observation Allocation

### v1

@@ 5.1.1. Satellite Constellations
We evaluate two constellations modeled after operational low Earth orbit constellations Planet (2023). The Planet constellation is modeled on the Dove constellation from Planet Labs. This constellation is composed of two near sun-synchronous orbital planes at $95^{\circ}$ inclinations each composed of $95$ satellites with an additional two orbital planes at $52^{\circ}$ inclinations each with $5$ satellites. The Walker constellation is motivated by the Skysat constellation from Planet Labs. This constellation has $6$ orbital planes with $14$ satellites each at an $88^{\circ}$ inclination and an overlay of $2$ orbital planes at a $51.6^{\circ}$ inclination with $12$ satellites. Each satellite has a memory capacity of $125$ GB and a single sensor that can slew to $60^{\circ}$ and $45^{\circ}$ off-nadir for the Planet and Walker constellations respectively. Figure 2 depicts these two constellations.
@@ 5.1.4. Hyperparameters and Execution Environment
The stochastic update of NSS and DSA relies on the probability, $P_{u}$ that determines when an agent should unassign from a task. We fix this value at $0.7$, as published in previous work Parjan and Chien (2023); Zilberstein et al. (2025b). The $maxIters$ is set to $20$. For the GND heuristic, we set all hyperparameters as described in previous work, and specifically use the GND(2) heuristic Zilberstein et al. (2025b). We use a random seed of $2005$ for the initial scenario generation. When evaluating $N$ scenarios, we increment this seed for generation of each subsequent simulation. The random seed used in the repair procedure is initialized to $1$. The NSS and DSA algorithms are seeded with a value of $1234$, and the random seed for GND is set to $2$. The random scheduling algorithm uses a seed of $2023$. All experiments are executed using Java 19 on a MacBook Pro 16 laptop with an M2 Max processor (12-core CPU and 38-core GPU) and 64 GB of RAM.
@@ 6. Conclusion
D-NSS generalizes to many DDCOP problems. While we leverage the GND heuristic, the D-NSS framework can accommodate alternative heuristics, including learned or adaptive ones, enabling broader applicability across DDCOP domains. Some examples include multi-agent path finding Salzman and Stern (2020), mobile sensor teams Pertzovsky et al. (2024), and UAV coordination Pujol-Gonzalez et al. (2013).
In future work, we can extend D-NSS to reason about future dynamics. We can construct or learn heuristics that prioritize tasks that are more likely to yield utility when executed. Alternatively, we can extend DCOSP to be a Proactive DCOP (PDDCOP) Hoang et al. (2022) where we have priors on problem dynamics, and agents can generate robust offline solutions. However, PDDCOPs require a model of problem dynamics which may not be obtainable in practice.
DCOSP solutions will be demonstrated in operational scenarios beginning in 2026. One example is NASA’s FAME demonstration Chien et al. (2025a) which centers around a federated observation system Zilberstein et al. (2025a). The FAME demonstration consists of more than 60 participating Earth-observing spacecraft that will coordinate their measurements to optimize observation completion across dynamic scenarios. Through onboard control, satellites will be able to react faster and with more precision to time-sensitive events such as natural disasters. These observations will provide novel measurements of scientific processes and crucial, up-to-date information for human operators. This work has laid the foundation for dynamic, large-scale, distributed onboard scheduling of Earth-observing satellites.
@@
Portions of the research were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). Government sponsorship acknowledged.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE2140739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ 5.1.1. Satellite Constellations
We evaluate two constellations modeled after operational low Earth orbit constellations planet. The Planet constellation is modeled on the Dove constellation from Planet Labs. This constellation is composed of two near sun-synchronous orbital planes at $95^{\circ}$ inclinations each composed of $95$ satellites with an additional two orbital planes at $52^{\circ}$ inclinations each with $5$ satellites. The Walker constellation is motivated by the Skysat constellation from Planet Labs. This constellation has $6$ orbital planes with $14$ satellites each at an $88^{\circ}$ inclination and an overlay of $2$ orbital planes at a $51.6^{\circ}$ inclination with $12$ satellites. Each satellite has a memory capacity of $125$ GB and a single sensor that can slew to $60^{\circ}$ and $45^{\circ}$ off-nadir for the Planet and Walker constellations respectively. Figure 2 depicts these two constellations.
@@ 5.1.4. Hyperparameters and Execution Environment
The stochastic update of NSS and DSA relies on the probability, $P_{u}$ that determines when an agent should unassign from a task. We fix this value at $0.7$, as published in previous work shreya; zilberstein2025decentralized. The $maxIters$ is set to $20$. For the GND heuristic, we set all hyperparameters as described in previous work, and specifically use the GND(2) heuristic zilberstein2025decentralized. We use a random seed of $2005$ for the initial scenario generation. When evaluating $N$ scenarios, we increment this seed for generation of each subsequent simulation. The random seed used in the repair procedure is initialized to $1$. The NSS and DSA algorithms are seeded with a value of $1234$, and the random seed for GND is set to $2$. The random scheduling algorithm uses a seed of $2023$. All experiments are executed using Java 19 on a MacBook Pro 16 laptop with an M2 Max processor (12-core CPU and 38-core GPU) and 64 GB of RAM.
@@ 6. Conclusion
D-NSS generalizes to many DDCOP problems. While we leverage the GND heuristic, the D-NSS framework can accommodate alternative heuristics, including learned or adaptive ones, enabling broader applicability across DDCOP domains. Some examples include multi-agent path finding mapf, mobile sensor teams ca-dcop-mst, and UAV coordination uavs-coord.
In future work, we can extend D-NSS to reason about future dynamics. We can construct or learn heuristics that prioritize tasks that are more likely to yield utility when executed. Alternatively, we can extend DCOSP to be a Proactive DCOP (PDDCOP) pddcop where we have priors on problem dynamics, and agents can generate robust offline solutions. However, PDDCOPs require a model of problem dynamics which may not be obtainable in practice.
DCOSP solutions will be demonstrated in operational scenarios beginning in 2026. One example is NASA’s FAME demonstration fame-spaceops-2025 which centers around a federated observation system zilberstein-spaceops-2025. The FAME demonstration consists of more than 60 participating Earth-observing spacecraft that will coordinate their measurements to optimize observation completion across dynamic scenarios. Through onboard control, satellites will be able to react faster and with more precision to time-sensitive events such as natural disasters. These observations will provide novel measurements of scientific processes and crucial, up-to-date information for human operators. This work has laid the foundation for dynamic, large-scale, distributed onboard scheduling of Earth-observing satellites.
@@
Portions of the research were carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). Government sponsorship acknowledged. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No DGE2140739. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.